Skip to main content

Problem 102: Amaker of specialized instruments receives shipments of 24 circuit boards. Suppose one shipment contains 4 that are defective. An engineer selects a random sample of size 4. What are the probabilities that the sample will contain
a. 0 defective circuit boards?
b. 1 defective circuit board ?
c. 2 or more defective circuit boards?

Problem 102: Amaker of specialized instruments receives shipments of 24 circuit boards. Suppose one shipment contains 4 that are defective. An engineer selects a random sample of size 4. What are the probabilities that the sample will contain
a. 0 defective circuit boards?
b. 1 defective circuit board ?
c. 2 or more defective circuit boards?


Solution:


a. 0 defective circuit boards
Given that $x=0,n=4,a=4, N=24$
Hypergeometric distribution, $h(x;n,a,N)=\frac{\left(\begin{array}{c}a\\ x\end{array}\right)\left(\begin{array}{c}N-a\\ n-x\end{array}\right)}{\left(\begin{array}{c}N\\ n\end{array}\right)}$
$P(x=0) = h(0;4,4,24)=\frac{\left(\begin{array}{c}4\\ 0\end{array}\right)\left(\begin{array}{c}24-4\\ 4-0\end{array}\right)}{\left(\begin{array}{c}24\\ 4\end{array}\right)} = 0.456.$

b. 1 defective circuit board
Given that $x=1,n=4,a=4, N=24$
Hypergeometric distribution, $h(x;n,a,N)=\frac{\left(\begin{array}{c}a\\ x\end{array}\right)\left(\begin{array}{c}N-a\\ n-x\end{array}\right)}{\left(\begin{array}{c}N\\ n\end{array}\right)}$
$P(x=1) = h(1;4,4,24)=\frac{\left(\begin{array}{c}4\\ 1\end{array}\right)\left(\begin{array}{c}24-4\\ 4-1\end{array}\right)}{\left(\begin{array}{c}24\\ 4\end{array}\right)} = 0.4291.$

c. 2 or more defective circuit boards
Given that $n=4,a=4, N=24$
Hypergeometric distribution, $h(x;n,a,N)=\frac{\left(\begin{array}{c}a\\ x\end{array}\right)\left(\begin{array}{c}N-a\\ n-x\end{array}\right)}{\left(\begin{array}{c}N\\ n\end{array}\right)}$
$P(x\ge 2) = 1-P(x<2) = 1-[P(x=0)+P(x=1)]$
$P(x=0) = h(0;4,4,24)=\frac{\left(\begin{array}{c}4\\ 0\end{array}\right)\left(\begin{array}{c}24-4\\ 4-0\end{array}\right)}{\left(\begin{array}{c}24\\ 4\end{array}\right)} = 0.456$
$P(x=1) = h(1;4,4,24)=\frac{\left(\begin{array}{c}4\\ 1\end{array}\right)\left(\begin{array}{c}24-4\\ 4-1\end{array}\right)}{\left(\begin{array}{c}24\\ 4\end{array}\right)} = 0.4291$
$P(x\ge 2) = 1-[P(x=0)+P(x=1)] = 1-(0.456+0.4291) = 1- 0.88509 = 0.1149$

Comments

Popular posts from this blog

Problem 105: If a bank receives on the average α = 6 bad checks per day, what are the probabilities that it will receive (a) 4 bad checks on any given day?
(b) 10 bad checks over any 2 consecutive days?

Problem 105: If a bank receives on the average α = 6 bad checks per day, what are the probabilities that it will receive (a) 4 bad checks on any given day? (b) 10 bad checks over any 2 consecutive days? Solution: Poisson distribution, $f(x;\lambda )=\frac{{\lambda }^{x}{e}^{-\lambda }}{x!}$ a. 4 bad checks on any given day Given that $x=4, \lambda = \alphaT = 6\dot 1 = 6$ $f(4;6 )=\frac{{6 }^{4}{e}^{-6 }}{4!} = 0.134$ b. 10 bad checks over any 2 consecutive days Given that $x=10, \lambda = \alphaT = 6\dot 2 = 12$ $f(10;12 )=\frac{{12 }^{10}{e}^{-12 }}{10!} = 0.104837$

Problem 101: Among the 300 employees of a company, 240 are union members, while the others are not, If 8 of the employees are chosen to serve on the administrative committee, find the probability that 5 of them will be union member while the others are not.
a. the formula for the hypergeometric distribution;
b. the formula for the binomial distribution as an approximation.

Problem 101: Among the 300 employees of a company, 240 are union members, while the others are not, If 8 of the employees are chosen to serve on the administrative committee, find the probability that 5 of them will be union member while the others are not. a. the formula for the hypergeometric distribution; b. the formula for the binomial distribution as an approximation. Solution: a. the formula for the hypergeometric distribution Given that $x=5,n=8,a=240, N=300$ Hypergeometric distribution, $h(x;n,a,N)=\frac{\left(\begin{array}{c}a\\ x\end{array}\right)\left(\begin{array}{c}N-a\\ n-x\end{array}\right)}{\left(\begin{array}{c}N\\ n\end{array}\right)}$ $P(x=5) = h(5;8,240,300)=\frac{\left(\begin{array}{c}240\\ 5\end{array}\right)\left(\begin{array}{c}300-240\\ 8-5\end{array}\right)}{\left(\begin{array}{c}300\\ 8\end{array}\right)} = 0.1470.$ b. the formula for the binomial distribution as an approximation? Given that $x=5,n=8,p = \frac{240}{300}$ $P(x= 5) =...

Problem 91: Suppose that a lot of 5000 electrical fuses contains 5% defectives. If a sample of 5 fuses is tested, find the probability of observing at least one defective.

Problem 91: Suppose that a lot of 5000 electrical fuses contains 5% defectives. If a sample of 5 fuses is tested, find the probability of observing at least one defective. Solution: Given that $n=5,$ and $p=0.05$ at least one defective; $P(x\ge 1) = 1-P(x=0)$ $b(x;n,p)=\left(\begin{array}{c}n\\ x\end{array}\right){p}^{x}(1-p)^{n-x}$ $P(x\ge 1) = 1 - b(0;5,0.05) = 1 - \left(\begin{array}{c}5\\ 0\end{array}\right){0.05}^{0}(1-0.05)^{5-0} = 1-(0.95)^5 = 0.226.$